OpenStudy (anonymous):

Find all the real-number roots of the equation. Give an exact expression for the root and also a calculator approximation rounded to three decimal places. (Enter your answers as comma-separated lists.) log9(x^2 + x) = 0.5

OpenStudy (anonymous):

log9(x^2 + x) = 0.5 x^2 +x = 9^(1/2) = 3 x^2 +x - 3 = 0 Delta = 1 + 12 = 13 x = -1 +/- sqrt (13)/ 2

OpenStudy (anonymous):

You can plug the numbers into your calculator now!